On a given 530 mile trip an SUV averaged 65mph during the first part and 35 mph during the second part. The total time of the trip was 10 hours. In how much time is each part of the trip?
Answer by TimothyLamb(4379)

average speed for entire trip:
530 / 10 = 53 mph
---
s = d / t
d = s * t
---
a = miles during first part
b = miles during second part
---
miles:
a + b = 530
d = s * t
a = 65x miles
b = 35y miles
---
x = hours during first part
y = hours during second part
---
hours:
x + y = 10
x = 10 - y
---
a + b = 530
65x + 35y = 530
65(10 - y) + 35y = 530
650 - 65y + 35y = 530
30y = 120
---
y = 4 hours
x = 6 hours
---
check:
a = 65(6) = 390 miles
b = 35(4) = 140 miles
a + b = 530
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php